# Hints for a Geon from Causal Dynamic Triangulations

### v1

@@ II Technical setup
For the creation of the configurations we use the code of [21]. For technical details of CDT simulations and notations, we refer to [5]. For the exploratory simulations at hand, we employed a single simulation point at the bare parameters $\Delta=0.6$ and $\kappa_{0}=2.2$, which exhibits a de Sitter-like space time [10], and thus a (classical) positive cosmological constant. Moreover, the lattice spacing, i. e. the (spatial) length of the simplices used to triangulate the space-time, has been estimated to be about $\approx 2.1$ Planck lengths [22], which we will take here at face value. We performed simulations on $N_{t}\times N_{s}=60\times 80$k, $80\times 160$k, and $80\times 320$k, with $N_{t}$ the temporal extent of the foliated space-time (i.e. the number of time slices, or spacing between timelike distances) and $N_{s}$ the number of simplices, which connect the time slices. Since we did not find any statistical significant dependence on the volume, we will show in the following only results from the largest volume [20]. We denote quantities at fixed foliation a time-slice, and everything on or in between two neighboring foliations a fat time-slice. In total, we calculated for each case 5000 (2000 for $N_{s}=320$k) configurations, following standard procedures for thermalization and decorrelation [23]. We choose spatially spherical boundary conditions at fixed foliation and periodic boundary condition in foliation direction. We note that the effective obtained geometry is found [22, 24] to be four-dimensional spherical, and thus (Euclidean) de Sitter.
@@
In order to extract the spatial distances, we could directly use the simplices’ spatial edges with a fixed distance. However, it turns out that this is not the best spatial discretization. Following [18] we employ a dual triangulation. For that, the center of neighboring simplices are used to define a discrete distance element. Though the unique association with a fixed time-slice is lost in this case, the pseudo-time-like distance between two consecutive time-slices is expected to become infinitesimal in the continuum limit [5, 10]. We therefore neglect this effect rather than introducing a fractional time as was done in [18]. Instead, we consider all centers of simplices in what we call a “fat” foliation slice to be at the same effective foliation count. While we do not see any obvious effects related to this approximation, this will be resolved in more detail in future work.
Distance measures between non-neighboring points themselves are now non-trivial. Following [5, 25, 6], we take as a space-like distance measure the geodesic distance $d$ between dual vertices on a fixed fat foliation slice. This is obtained by employing a Dijkstra algorithm, which is run on the triangulation restricted to a certain fat foliation slice. Since the determination using all possible pairs is, of course, prohibitively expensive, we have uniformly sampled 100 (50 for $N_{s}=80$k) pairs per fixed fat foliation slice for a stochastic estimator [26]. This is similar to the discussion in [18], in particular their remarks in Appendix B, which we also took into account.
@@
$Q(\epsilon)$ $=$ $\frac{q}{\delta}=\frac{1}{N(p,\delta)N(p^{\prime},\delta)}\sum_{\begin{subarray}{c}d(p,p^{\prime})=\epsilon\\ d(q,p)\leq\delta,d(q^{\prime},p^{\prime})\leq\delta\end{subarray}}d(q,q^{\prime})$
@@
where the sums run over all points of the dual triangulation, i.e. the centers of simplices. As $Q$ does no longer depend on $\epsilon$ explicitly, it is possible to identify $p$ and $p^{\prime}$ by setting $\epsilon=0$, thus yielding a quantity measuring the distance between pairwise points in a sphere with respect to each other [18]. Thus, $Q$ is a measure of such distances, and triangulations with different local structure will in general yield locally different values of $Q$.
If the theory has a smooth continuum limit to a Riemannian manifold then the limit $\delta\to 0$ yields [27]
$Q=q_{1}+\delta^{2}(-q_{2}vRv-q_{3}R+{\cal O}(\delta))$ (1)
where the $q_{i}$ are positive, real constants, $R$ is the standard, continuum, curvature scalar, and $vRv$ is the Ricci tensor projected on orthonormal tangent vectors $v$ at $p$. Thus, $Q$ is in this limit a measure for curvature. Indeed, this was confirmed to be a good approximation of the functional dependency on $\delta$ in [27] for CDT simulations. We confirmed this in the present work.
@@ III Defining a geon correlation function
It is well established [10, 24] that a single configuration, and even the average, in CDT simulations at the given simulation point do not resemble a flat Minkowski manifold, but a(n Euclidean) de Sitter space. This implies that it is impossible to speak about locations across configurations. In addition, the number of foliations is usually not very large compared to the number of foliations where the de Sitter space is almost static. It is a priori not clear, how the global space-time characteristics would affect the propagation of a gravitational state. We therefore opt for another approach.
@@ IV Results
Figure 3: The normalized correlation function for $O=\Delta Q$ with fits of type $a+b\exp(-ms)$.
@@
The shape of the correlators remains roughly the same, and gets only more compressed with increasing cosmological time, and thus decreasing spatial extent of the universe. However, as Fig. 4 shows, the mass parameter, determining the exponential decay, is relatively stable, except during the phase of rapid expansion, where it quickly rises. It remains, however, approximately the same for both operators. This suggest that during the expansion of the Universe in our CDT simulations, the properties of geons changes drastically, but are more or less constant at times of constant extension. This pattern is the same for all volumes, and we observe no statistically significant volume dependence. The masses, averaged over $120\leq\tau<$ 12, are found to be 0.18(1), 0.16(1), 0.17(3) for the $Q$ correlator and 0.14(1), 0.17(1), 0.15(3) for the $Q^{2}$ correlator, for $N_{s}$ being 80k, 160k, and 320k, respectively. It would correspond to a mass of $\sim 0.09M_{\text{Planck}}$.
@@ Acknowledgments
We are grateful to Dániel Németh for support, a critical reading of this manuscript, and him and all of the authors of [21] for providing us with the code from [21]. We are grateful to him and Renate Loll for discussions.

### v2

@@ II Technical setup
For the creation of the configurations we use the code of [21]. For technical details of CDT simulations and notations, we refer to [5]. For the exploratory simulations at hand, we employed a single simulation point at the bare parameters $\Delta=0.6$ and $\kappa_{0}=2.2$, which exhibits a de Sitter-like space time [10], and thus a (classical) positive cosmological constant. Moreover, the lattice spacing, i. e. the (spatial) length of the simplices used to triangulate the space-time, has been estimated to be about $\approx 2.1$ Planck lengths [22], which we will take here at face value. We performed simulations on $N_{t}\times N_{\text{simp}}=60\times 80$k, $80\times 160$k, and $80\times 320$k, with $N_{t}$ the temporal extent of the foliated space-time (i.e. the number of time slices, or spacing between timelike distances) and $N_{\text{simp}}$ the number of (4,1)-simplices, which connect the time slices. Since we did not find any statistical significant dependence on the volume, we will show in the following only results from the largest volume [20]. We denote quantities at fixed foliation a time-slice, and everything on or in between two neighboring foliations a fat time-slice. In total, we calculated for each case 5000 (2000 for $N_{\text{simp}}=320$k) configurations, following standard procedures for thermalization and decorrelation [23]. We choose spatially spherical boundary conditions at fixed foliation and periodic boundary condition in foliation direction. We note that the effective obtained geometry is found [22, 24] to be four-dimensional spherical, and thus (Euclidean) de Sitter.
@@
In order to extract the spatial distances, we could directly use the simplices’ spatial edges with a fixed distance. However, it turns out that this is not the best spatial discretization. Following [18] we employ a dual triangulation. For that, the center of neighboring simplices are used to define a discrete distance element. Though the unique association with a fixed time-slice is lost in this case, the pseudo-time-like distance between two consecutive time-slices is expected to become infinitesimal in the continuum limit [5, 10]. We therefore neglect this effect rather than introducing a fractional time as was done in [25, 24]. Instead, we consider all centers of simplices in what we call a “fat” foliation slice to be at the same effective foliation count. While we do not see any obvious effects related to this approximation, this will be resolved in more detail in future work.
Distance measures between non-neighboring points themselves are now non-trivial. Following [5, 26, 6], we take as a space-like distance measure the geodesic distance $d$ between dual vertices on a fixed fat foliation slice. This is obtained by employing a Dijkstra algorithm, which is run on the triangulation restricted to a certain fat foliation slice. Since the determination using all possible pairs is, of course, prohibitively expensive, we have uniformly sampled 100 (50 for $N_{\text{simp}}=80$k) pairs per fixed fat foliation slice for a stochastic estimator [27]. This is similar to the discussion in [18], in particular their remarks in Appendix B, which we also took into account.
@@
$Q(\epsilon)$ $=$ $\frac{q}{\delta}=\frac{1}{N(p,\delta)N(p^{\prime},\delta)}\sum_{\begin{subarray}{c}d(p,p^{\prime})=\epsilon\\ d(q,p)=\delta,d(q^{\prime},p^{\prime})=\delta\end{subarray}}d(q,q^{\prime})$
@@
where the sums run over all points of the dual triangulation, i.e. the centers of simplices. As $Q$ does no longer depend on $\epsilon$ explicitly but only through the summation, it is possible to identify $p$ and $p^{\prime}$ by setting $\epsilon=0$, thus yielding a quantity measuring the distance between pairwise points in a sphere with respect to each other [18]. Thus, $Q$ is a measure of such distances, and triangulations with different local structure will in general yield locally different values of $Q$. In practice, we fixed $\delta=6$ for this work. However, when varying $\delta$ in some range around this value, we did not see any significant change [20].
If the theory has a smooth continuum limit to a Riemannian manifold then the limit $\delta\to 0$ yields [28, 18]
$Q=q_{1}-\delta^{2}(q_{3}R+{\cal O}(\delta))$ (1)
where the $q_{i}$ are positive, real constants, $R$ is the standard, continuum, curvature scalar. Thus, $Q$ is in this limit a measure for curvature. Indeed, this was confirmed to be a good approximation of the functional dependency on $\delta$ in [28] for CDT simulations. We confirmed this in the present work.
@@ III Defining a geon correlation function
It is well established [10, 24] that a single configuration, and even the average, in CDT simulations at the given simulation point do not resemble a flat Minkowski manifold, but a(n Euclidean) de Sitter space. This implies that it is impossible to speak about locations across configurations. In addition, the number of time-slices is usually not very large compared to the number of time-slices where the de Sitter space is almost static. It is a priori not clear, how the global space-time characteristics would affect the propagation of a gravitational state. We therefore opt for another approach.
@@ IV Results
Figure 3: The normalized correlation function for $O=\Delta Q$ with fits of type $a+b\exp(-ms)$ and their ±$\pm 1\sigma$ error band.
@@
The shape of the correlators remains roughly the same, and gets only more compressed with increasing cosmological time, and thus decreasing spatial extent of the universe. However, as Fig. 4 shows, the mass parameter, determining the exponential decay, is relatively stable, except during the phase of rapid expansion, where it quickly rises. It remains, however, approximately the same for both operators. This suggest that during the expansion of the Universe in our CDT simulations, the properties of geons changes drastically, but are more or less constant at times of constant extension. This pattern is the same for all volumes, and we observe no statistically significant volume dependence. The masses, averaged over $120\leq\tau<$ 12, are found to be 0.18(1), 0.16(1), 0.17(3) for the $Q$ correlator and 0.14(1), 0.17(1), 0.15(3) for the $Q^{2}$ correlator, for $N_{\text{simp}}$ being 80k, 160k, and 320k, respectively. It would correspond to a mass of $\sim 0.09M_{\text{Planck}}$.
@@ Acknowledgments
We are grateful to Dániel Németh for support, and him and all of the authors of [21] for providing us with the code from [21]. We are grateful to him and Renate Loll for discussions. We are grateful to Renate Loll, Dániel Németh, and Jesse van der Duin for useful comments on the manuscript. Simulations have been performed on the GSC at Graz University.